# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ F.2.2   Helpman-Book1998 and Redding-Sturm-AER2008 model
Helpman-Book1998 removed the A-sector in the Krugman model and assumed that all workers are mobile. Instead of the A-sector, the model introduces the housing (abbreviated as H) sector. Each region $i$ is endowed with a fixed stock $a_{i}$ of housing. Workers’ preference is Cobb–Douglas of M-sector CES aggregate $C_{i}^{\mathrm{M}}$ and H-sector goods $C_{i}^{\mathrm{H}}$, $u_{i}=(C_{i}^{\mathrm{M}})^{\mu}(C_{i}^{\mathrm{H}})^{\gamma}$, where $\mu\in(0,1)$ is the expenditure share of the former and $\gamma=1-\mu\in(0,1)$ is that for the latter. There are two variants for assumptions on how housing stocks are owned: public landownership (abbreviated as PL) and local landownership (LL). The original formulation by Helpman-Book1998 supposes housing stocks are equally owned by all workers (i.e., PL). The income of a worker in region $i$ is the sum of the wage and an equal dividend $r>0$ of the total rental revenue in the economy. However, ottaviano2002agglomeration, Murata-Thisse-JUE2005, and Redding-Sturm-AER2008 assumed that housing stocks are locally owned (i.e., LL). The income of a worker in region $i$ is the sum of the wage and an equal dividend of rental revenue in each region. In fact, the model by Redding-Sturm-AER2008 is the LL version of the Helpman model.
@@
[LL] $e_{i}=w_{i}x_{i}.$ (F.24)
Also, the market wage is given as the solution for Eq. F.5. For the PL case, we set $r=1$ for normalization. For the LL case, $\bm{w}(\bm{x})$ is uniquely given up to normalization. The indirect utility function is, with $\Delta_{i}\equiv\sum_{j\in\mathcal{I}}x_{j}w_{j}^{1-\sigma}\phi_{ji}$ and $r>0$,
@@
[LL] $v_{i}=\left(\frac{x_{i}}{a_{i}}\right)^{-\gamma}w_{i}^{\mu}\Delta_{i}^{\frac{\mu}{\sigma-1}}.$ (F.26)
@@
[…]
and $\mathbf{M}$ is defined by Eq. F.6. For the PL case, we obtain $\mathbf{V}=\Omega(\mathbf{D})$ with
@@
where we compute $\mathbf{W}_{x}$ from Eq. F.7 with $\zeta=\frac{\mu(\bar{w}+1)}{\bar{w}}$, $\epsilon_{x}=\frac{\bar{w}+1}{\bar{w}}$, $\epsilon_{w}=\mu$; we note that $\frac{\bar{w}}{\bar{w}+1}=\mu$ under our normalization. Thus, for the PL case, we can choose $\Omega^{\sharp}$ and $\Omega^{\flat}$ that satisfy $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$ as follows:
@@
Remark 8.
As in the Krugman model (Remark 5), one can identify the correspondence between the coefficients in $\Omega$ Eq. F.31 and those in $\Omega^{\sharp}$ Eq. F.30. For example, local dispersion force $-\gamma$ in Eq. F.30 affects not only constant term in $\Omega^{\sharp}(\cdot)$ but also the first- and the second-order terms of $\Theta$. Other forces can be understood similarly to the Krugman model (Remark 5). ∎
For the LL case, $\mathbf{W}_{x}$ is given in Example 4 and we obtain
@@
Definition 2 classifies canonical models based on the spatial scale of the “effective” dispersion forces. The Helpman model with public landownership Eq. F.31 has global dispersion forces because $c_{2}<0$ under Assumption E. However, unlike the Krugman model, the global dispersion forces in the Helpman model are not “effective” in the sense that, under any admissible values of $\mu$ and $\sigma$, this force does not stabilize the uniform distribution for any level of transport costs. If we drop the local dispersion force $c_{0}<0$ from $\Omega^{\sharp}$, we have $c_{1}\Theta+c_{2}\Theta^{2}>0$ for all $\Theta$ and $\bar{\bm{x}}$ is always unstable. Thus, the only dispersion force in the Helpman model that can stabilize $\bar{\bm{x}}$ is its local dispersion force. ∎
@@
Equilibrium is unique when $\gamma\sigma=(1-\mu)\sigma>1$ (Redding-Sturm-AER2008). For both PL and LL, it implies that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(0,1)$. ∎
@@
The payoff function is given by Eq. F.29 with $\Delta_{i}=\sum_{k\in\mathcal{I}}x_{k}a_{k}w_{k}^{1-\sigma}\phi_{ki}$.
From Example 5, $\mathbf{S}_{a}=-\frac{\bar{w}\bar{x}}{\bar{a}}\left(\mathbf{I}-\mathbf{D}\right)\left(\mathbf{I}+\mathbf{D}\right)$ as $\bar{e}=\bar{w}\bar{x}$. Also, we have $\tilde{\mathbf{V}}_{w}=\frac{\bar{v}}{\bar{w}}\mu(\mathbf{I}-\mathbf{D})$, $\tilde{\mathbf{V}}_{a}=\frac{\bar{v}}{\bar{a}}\frac{\mu}{\sigma-1}\mathbf{D}$, and $\mathbf{S}_{w}=\sigma\bar{x}\Omega^{\flat}(\mathbf{D})$. As $\mathbf{V}_{a}=\tilde{\mathbf{V}}_{a}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{a}$ and $\mathbf{A}=\frac{\bar{a}}{\bar{v}}\mathbf{V}_{a}=\Omega^{\natural}(\mathbf{D})$, we compute
@@
where $\Omega^{\sharp}(\Theta)$ is that for the LL case Eq. F.35. We have $\delta^{\prime}(\Theta)>0$ for all $\Theta$ whenever $(1-\mu)\sigma>1$. ∎
@@ F.2.3   Pfluger-Suedekum-JUE2008 model
The Pflüger–Südekum model builds on Pfluger-RSUE2003 and introduces the housing sector (again denoted by H). In this model, a quasi-linear upper-tier utility is assumed: $u_{i}=C_{i}^{\mathrm{A}}+\mu\log C_{i}^{\mathrm{M}}+\gamma\log C_{i}^{\mathrm{H}}$. The production cost for a firm in $i\in\mathcal{I}$ is $\alpha w_{i}+\beta q$. Then, $\bm{w}$ is given as follows:
@@
where $\Delta_{i}=\sum_{j\in\mathcal{I}}\phi_{ji}x_{j}$, and $l_{i}$ and $a_{i}$ denote the mass of immobile workers and amount of housing stock in region $i$, respectively. The nominal wage in region $i$ is given by Eq. F.39. Let $l_{i}=l$ and $a_{i}=a$ for all $i$. Then, we see that $\mathbf{V}=\frac{1}{\bar{v}}\Omega^{\sharp}(\mathbf{D})$ with
@@ F.2.4   Allen-Arkolakis-QJE2014 model
The AA model is a perfectly competitive Armington-IMF1969-based framework with positive and negative local externalities. We consider a discrete-space version of the model and abstract away all exogenous differences across regions. In the model, productivity of a location is proportional to $x_{i}^{\alpha}$ with $\alpha>0$, representing positive externalities. The market equilibrium condition is
@@
With market wage $\bm{w}$, we have $v_{i}(\bm{x})=x_{i}^{\beta}w_{i}\Delta_{i}^{\frac{1}{\sigma-1}}$ with $\Delta_{i}\equiv\sum_{k\in\mathcal{I}}w_{k}^{1-\sigma}x_{k}^{\alpha(\sigma-1)}\phi_{ki}$; $x_{i}^{\beta}$ ($\beta<0$) represents negative externalities from congestion. We have $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$ with
@@
It is also instructive to consider the case with no externalities ($\alpha=0$ and $\beta=0$). Then, the model reduces to the Armington-IMF1969 trade model, and we have $\omega=-\frac{1}{\sigma(\sigma+(\sigma-1)\Theta)}<0$. Thus, the intrinsic working of general equilibrium effects induced by love for variety in the Armington model is in creating a dispersion force, and it is global in the sense that the force depends on $\Theta$. However, in the context of net agglomeration incentive at the symmetry with nonzero spillovers, this force is mainly related to the denominator $\Omega^{\flat}$, and $\Omega^{\sharp}$ succinctly summarizes the net trade-off between agglomeration and dispersion forces that govern the stability of $\bar{\bm{x}}$.
Figure F.1 provides the classification of spatial patterns and their stability under Assumption C. Figure F.1 can be seen as a refinement of Figure I in Allen-Arkolakis-QJE2014 under Assumption C. A sufficient condition for the equilibrium uniqueness is $\alpha+\beta\leq 0$ (Range 3), which means there is net local dispersion force and $\Omega^{\sharp}(\Theta)<0$ for all $\Theta$. In Range 3, $\bar{\bm{x}}$ is the only equilibrium.
Remark 13.
In Fig. C.3, we set $(\alpha,\beta,\sigma)=(0.5,-0.3,6)$. ∎
@@ F.2.6   Kucheryavyy-etal-JIE2024
As seen, the model is Type LG in its most general form. Importantly, the case $(\alpha,\zeta,\mu)=(1,1,\beta)$ corresponds to a Krugman-type model, as we observe Eq. F.46 reduces to a function of the form $c_{1}\Theta-c_{2}\Theta^{2}$. Also, $\alpha<1$ and $\zeta=1$ corresponds to the Allen–Arkolakis framework, for which we confirm $-(1-\alpha)<0$ represents a local dispersion force. It is noted that $\Omega^{\sharp}(1)=(1-\mu)(1-\beta)$. If $\mu<1$, then $\Omega^{\sharp}(1)>0$, so that $\Omega^{\sharp}$ must have one and one zero for $\Theta\in(0,1)$, so that $\bar{\bm{x}}$ is stable only for low transport costs.
References for Supplementary Materials

### v8

@@ F.2.2  Helpman-Book1998 and Redding-Sturm-AER2008 model
Helpman-Book1998 removed the A-sector in the Krugman model and assumed that all workers are mobile, and introduced the housing sector (abbreviated as H). Each region $i$ is endowed with a fixed stock $a_{i}$ of housing. Workers’ preference is Cobb–Douglas of M-sector CES aggregate $C_{i}^{\mathrm{M}}$ and H-sector goods $C_{i}^{\mathrm{H}}$, $u_{i}=(C_{i}^{\mathrm{M}})^{\mu}(C_{i}^{\mathrm{H}})^{\gamma}$, where $\mu\in(0,1)$ is the expenditure share of the former and $\gamma=1-\mu\in(0,1)$ is that for the latter.
There are two variants for assumptions on how housing stocks are owned: public landownership (PL) and local landownership (LL). Helpman-Book1998 supposes PL in which housing stocks are equally owned by all workers; the income of a worker in region $i$ is the sum of the wage and an equal dividend $r>0$ of the total rental revenue in the economy. However, ottaviano2002agglomeration, Murata-Thisse-JUE2005, and Redding-Sturm-AER2008 assumed that housing stocks are locally owned (i.e., LL). The income of a worker in region $i$ is the sum of the wage and an equal dividend of rental revenue in each region. In fact, the model by Redding-Sturm-AER2008 is the LL version of the Helpman model.
@@
[$L$] $e_{i}=w_{i}x_{i}.$ (F.24)
Also, the market wage is given as the solution for Eq.˜F.5. For the PL case, we set $r=1$ for normalization. For the LL case, $\bm{w}(\bm{x})$ is uniquely given up to normalization. The indirect utility function is, with $\Delta_{i}\equiv\sum_{j\in\mathcal{I}}x_{j}w_{j}^{1-\sigma}\phi_{ji}$ and $r>0$,
@@
[$L$] $v_{i}=\left(\frac{x_{i}}{a_{i}}\right)^{-\gamma}w_{i}^{\mu}\Delta_{i}^{\frac{\mu}{\sigma-1}}.$ (F.26)
@@
where $\text{[PL]}\quad\hat{\mathbf{V}}_{w}\equiv\mu\left(\operatornamewithlimits{diag}[\bm{w}+r\bm{1}]^{-1}-\mathbf{M}^{\top}\operatornamewithlimits{diag}[\bm{w}]^{-1}\right),$ (F.28)
$\text{[LL]}\quad\hat{\mathbf{V}}_{w}\equiv\mu\left(\mathbf{I}-\mathbf{M}^{\top}\right)\operatornamewithlimits{diag}[\bm{w}]^{-1},$ (F.29)
and $\mathbf{M}$ is defined by Eq.˜F.6. For the PL case, we obtain $\mathbf{V}=\Omega(\mathbf{D})$ with
@@
where we compute $\mathbf{W}_{x}$ from Eq.˜F.7 with $\zeta=\frac{\mu(\bar{w}+1)}{\bar{w}}$, $\epsilon_{x}=\frac{\bar{w}+1}{\bar{w}}$, $\epsilon_{w}=\mu$; we note that $\frac{\bar{w}}{\bar{w}+1}=\mu$ under our normalization. Thus, for the PL case, we can choose $\Omega^{\sharp}$ and $\Omega^{\flat}$ that satisfy $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$ as follows:
@@
For the LL case, $\mathbf{W}_{x}$ is given in Example˜6 and we obtain
@@
Definition˜3 classifies canonical models based on the spatial scale of the “effective” dispersion forces. The Helpman model with public landownership Eq.˜F.31 has global dispersion forces because $c_{2}<0$ under Assumption˜E. However, unlike the Krugman model, the global dispersion forces in the Helpman model are not “effective” in the sense that, under any admissible values of $\mu$ and $\sigma$, this force does not stabilize the uniform distribution for any level of transport costs. If we drop the local dispersion force $c_{0}<0$ from $\Omega^{\sharp}$, we have $c_{1}\Theta+c_{2}\Theta^{2}>0$ for all $\Theta$ and $\bar{\bm{x}}$ is always unstable. Thus, the only dispersion force in the Helpman model that can stabilize $\bar{\bm{x}}$ is its local dispersion force. ∎
@@
Equilibrium is unique when $\gamma\sigma=(1-\mu)\sigma>1$ (Redding-Sturm-AER2008). For both PL and LL, this condition implies that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(0,1)$. ∎
@@
The payoff function is given by Eq.˜F.29 with $\Delta_{i}=\sum_{k\in\mathcal{I}}x_{k}a_{k}w_{k}^{1-\sigma}\phi_{ki}$.
From Example˜7, $\mathbf{S}_{a}=-\frac{\bar{w}\bar{x}}{\bar{a}}\left(\mathbf{I}-\mathbf{D}\right)\left(\mathbf{I}+\mathbf{D}\right)$ as $\bar{e}=\bar{w}\bar{x}$. Also, we have $\tilde{\mathbf{V}}_{w}=\frac{\bar{v}}{\bar{w}}\mu(\mathbf{I}-\mathbf{D})$, $\tilde{\mathbf{V}}_{a}=\frac{\bar{v}}{\bar{a}}\frac{\mu}{\sigma-1}\mathbf{D}$, and $\mathbf{S}_{w}=\sigma\bar{x}\Omega^{\flat}(\mathbf{D})$. As $\mathbf{V}_{a}=\tilde{\mathbf{V}}_{a}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{a}$ and $\mathbf{A}=\frac{\bar{a}}{\bar{v}}\mathbf{V}_{a}=\Omega^{\natural}(\mathbf{D})$, we compute
@@
where $\Omega^{\sharp}(\Theta)$ is that for the LL case Eq.˜F.35. If $(1-\mu)\sigma>1$, $\delta^{\prime}(\Theta)>0$ for all $\Theta$. ∎
@@ F.2.3  Pfluger-Suedekum-JUE2008 model
The Pflüger–Südekum model builds on Pfluger-RSUE2003 and introduces the housing sector (denoted by H). A quasi-linear upper-tier utility is assumed: $u_{i}=C_{i}^{\mathrm{A}}+\mu\log C_{i}^{\mathrm{M}}+\gamma\log C_{i}^{\mathrm{H}}$. The production cost for a firm in $i\in\mathcal{I}$ is $\alpha w_{i}+\beta q$. Then, $\bm{w}$ is given as follows:
@@
where $\Delta_{i}=\sum_{j\in\mathcal{I}}\phi_{ji}x_{j}$, and $l_{i}$ and $a_{i}$ denote the mass of immobile workers and amount of housing stock in region $i$, respectively. The nominal wage in region $i$ is given by Eq.˜F.39. Let $l_{i}=l$ and $a_{i}=a$ for all $i$. Then, we see that $\mathbf{V}=\frac{1}{\bar{v}}\Omega^{\sharp}(\mathbf{D})$ with
@@ F.2.4  Allen-Arkolakis-QJE2014 model
The Allen–Arkolakis model is a perfectly competitive Armington-IMF1969-based framework with positive and negative local externalities. We abstract away all exogenous differences in regional fundamentals. The productivity of region $i$ is proportional to $x_{i}^{\alpha}$ with $\alpha>0$, representing positive externalities. The market equilibrium condition is
@@
With market wage $\bm{w}$, we have $v_{i}(\bm{x})=x_{i}^{-\beta}w_{i}\Delta_{i}^{\frac{1}{\sigma-1}}$ with $\Delta_{i}\equiv\sum_{k\in\mathcal{I}}w_{k}^{1-\sigma}x_{k}^{\alpha(\sigma-1)}\phi_{ki}$. With $\beta<0$, $x_{i}^{\beta}$ represents negative externalities from congestion. Here, we follow the original study in terms of the sign of $\beta$, while the main text uses $\beta$ as the magnitude of this externality to streamline exposition. We have $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$ with
@@
The case with no externalities ($\alpha=0$ and $\beta=0$) reduces to the Armington-IMF1969 framework, and $\omega=-\frac{1}{\sigma(\sigma+(\sigma-1)\Theta)}<0$. The intrinsic working of general equilibrium effects induced by love for variety in the Armington model is in creating a dispersion force, and it is in a sense “global” because it depends on $\Theta$. However, in the context of net agglomeration incentive at the symmetry with nonzero spillovers, this force is mainly related to the denominator $\Omega^{\flat}$, and $\Omega^{\sharp}$ summarizes the net trade-off between agglomeration and dispersion forces that govern the stability of $\bar{\bm{x}}$.
Figure˜F.1 classifies spatial patterns and their stability under Assumption˜C, which can be seen as a refinement of Figure I in Allen-Arkolakis-QJE2014 under Assumption˜C. A sufficient condition for the equilibrium uniqueness is $\alpha+\beta\leq 0$ (Range 3), which means there is net local dispersion force and $\Omega^{\sharp}(\Theta)<0$ for all $\Theta$. In Range 3, $\bar{\bm{x}}$ is the only equilibrium.
@@ F.2.6  Kucheryavyy-etal-JIE2024 model
As seen, the model is Type LG in its most general form. Importantly, the case $(\alpha,\zeta,\mu)=(1,1,\beta)$ corresponds to a Krugman-type model, as we observe Eq.˜F.46 reduces to a function of the form $c_{1}\Theta-c_{2}\Theta^{2}$. Also, $\alpha<1$ and $\zeta=1$ corresponds to the Allen–Arkolakis framework, for which we confirm $(1-\alpha)<0$ represents a local dispersion force. It is noted that $\Omega^{\sharp}(1)=(1-\mu)(1-\beta)$. If $\mu<1$, then $\Omega^{\sharp}(1)>0$, so that $\Omega^{\sharp}$ must have one and one zero for $\Theta\in(0,1)$, so that $\bar{\bm{x}}$ is stable only for low transport costs.